OpenStudy (anonymous):

What is the mass of an atom that has 4 protons, 5 neutrons and 4 electrons?

OpenStudy (anonymous):

\[mass of proton,m_p = 1.6726 \times 10^{-27} kg\] \[mass of neutron ,m_n = 1.6749 \times 10^{-27} kg\] \[mass of electron,m_e= 9.10 \times 10^{-31} kg\] mass of atom = \[4m_p +5m_n +4m_e \approx4m_p +5m_n \]

OpenStudy (shivaniits):

you could ignore the mass of electron as it relatively lighter than protons and neutrons..considering mass of protons and neutrons the mass would be 9 amu(atomic mass unit) or u could say that 9 gms of it would make a mole(\[6.023\times 10^{23}\] atoms or molecules of the substance....!!